Solve the following system of equations and provide a graphical representation of the solution.



y - x = 1



x2 + y2 = 13


1
Expert's answer
2022-06-01T08:52:36-0400

y=1+x

x2+(1+x)2=13

x2+1+2x+x2-13=0

2x2+2x-12=0

x2+x-6=0

D=1+24=25

x1=(-1+5)/2=2

x2=(-1-5)/2=-3

y1=1+2=3

y2=1-3=-2

Answer (2; 3), (-3; -2)